# What is Wrong with Perplexity for Long-context Language Modeling?

### v1

@@ 4 Experiments
In this section, we conduct real-world experiments to analyze the applicability of the proposed LongPPL and LongCE. For all the experiments, we use LongBench (Bai et al., 2023b), LongEval (Li et al., 2023a) and RULER (Hsieh et al., 2024) as the long-context benchmarks. LongBench mainly contains real-world tasks, including multi-document QA and summarization, while LongEval and RULER focus on synthetic tasks such as key-value retrieval and needle-in-a-haystack (NIAH). We report the average score on LongBench, the accuracy on the subtask “lines” of LongEval, and the score on RULER. For LongBench and RULER, we restrict the prompt length to 32k tokens. For LongEval, we use 1350 lines as the prompt, which is approximately 32k tokens.
@@ 4.1 LongPPL Metric
Experimental Setup. We calculate LongPPL on the GovReport dataset (Huang et al., 2021), which consists of long sequences from government reports. We sample 50 documents with the context length up to 32k tokens. We set the hyperparameters as $\alpha=2,\beta=-2,K=4096$. We use Qwen2-72B-Instruct (Yang et al., 2024), an open-source LLM with the context length of 128k tokens, as the discriminator model $\theta_{0}$ to select the key tokens.
@@
Figure 5: Correlation between the PPL-based metrics (LongPPL and PPL) on GovReport (Huang et al., 2021) and long-context benchmarks. LongPPL is calculated using Qwen2-72B-Instruct.
(a) Mistral Large 2
(b) Llama-3.1-8B
Figure 6: Correlation between LongBench score and LongPPL on GovReport calculated using Mistral Large 2 and Llama-3.1-8B.
LongPPL Correlates Well with Long-context Performance. In Figure 1(b) and 5, we demonstrate the correlation between LongPPL and long-context benchmarks on various long-context LLMs. We observe that LongPPL exhibits a very strong negative correlation with performance on long-context tasks across different models, with pearson correlation coefficients exceeding -0.8 for all three tasks. In contrast, perplexity hardly shows a correlation with the long-context tasks. This indicates that LongPPL is sufficiently capable of measuring a model’s long-context capabilities.
Table 1: The Pearson correlation between different perplexity measures and benchmark scores, which is lower the better (because we expect a lower perplexity indicates higher benchmark scores).
@@
LongPPL-hard (default) $I$ (Equation 4) -0.97 -0.85 -0.83
LongPPL is compatible with various evaluator models. To demonstrate that LongPPL is not restricted to a specific evaluator model, we additionally tested it with two other powerful long-context LLMs, Mistral Large 2, Llama-3.1-8B. As shown in Table 6, the LongPPL calculated using different models as the evaluator model also exhibits high correlation. We observe that even though the Llama-3.1 model we use is only 8B in size, it still obtains LongPPL with a strong correlation. This suggests that LongPPL is highly compatible with powerful long-context models, even with smaller parameter sizes.
Hard standard for key tokens is better than soft re-weighting standard. In Equation 4, we use an indicator function $I$ as the influence function. Instead, we have also tried to use the soft reweighting function $I_{\rm soft}$ used in LongCE (Equation 7) to calculate LongPPL. Its token matching strategy is detailed in Appendix A.1. In Table 1, we show that LongPPL with soft criteria has a weaker correlation with the long-context benchmarks compared to LongPPL, indicating that the soft reweighting influence function is suboptimal for LongPPL.
@@ 4.2 Fine-tune with LongCE loss
Experimental Setup. We primarily use Llama-2-7B (Touvron et al., 2023) as the base model to perform long-context finetuning. We also conduct experiments on Mistral-7B-v0.1 (Jiang et al., 2023) and Llama-2-13B. We use PG-19 (Rae et al., 2019), a book dataset sourced from a library, and Pile-arxiv (Gao et al., 2020), a dataset consisting of Arxiv papers, as the training dataset. The training sequences are organized to be the context length with 32k tokens. For the calculation of LongCE, we set $\gamma=5$ in Equation 7 and use the same sliding window approach as described in Section 4.1 to improve training efficiency. The context length of $\bm{s}_{i}$ is set to be $K=4096$. We fine-tune the base models with Entropy-aware Adjusted Base Frequency (EABF) (Zhang et al., 2024c) and Position Interpolation (PI) (Chen et al., 2023). Specifically, EABF applies a scaling mechanism to the attention and uses a higher base frequency for RoPE, while PI linearly downscales the position indices of the input tokens. These methods can significantly accelerate the convergence speed of long-context fine-tuning and have been widely adopted in many LLMs (Yang et al., 2024; Dubey et al., 2024; Chen et al., 2024a). Detailed training setups are available in Appendix A.2.
@@
LongCE Outperforms CE in Various Settings. As shown in Table 2, we present the long-context capabilities of models fine-tuned with LongCE loss and CE loss under different fine-tuning strategies and training datasets. We also test the effectiveness of LongCE using different base models in Table 3. We find that models fine-tuned with LongCE loss consistently outperform those fine-tuned with CE loss across nearly all settings. This suggests that the LongCE loss, with its re-weighting strategy based on long-context token importance, can be applied as a plug-and-play module which can effectively improve the model’s long-context performance. We also provide fine-grained results in Appendix B.3 to illustrate the specific task domains in which LongCE outperforms the baseline.
@@
Training Efficiency. In addition to the performance improvement brought by the LongCE loss, we also pay attention to the changes in training efficiency. In LongCE, we need an extra forward pass to calculate the probability under short context $P_{\theta}(x_{i}|\bm{s}_{i})$, which introduces additional computation costs. By using a sliding window technique (detailed in Appendix A.1), the computational overhead of training the model with LongCE is controlled to about 80% that of training with CE loss. In most settings shown in Table 2, the model trained with LongCE outperforms the model trained with CE, even when the latter is trained for twice as many steps. Additionally, we visualize in Figure 7 how the long-context performance of models fine-tuned with LongCE and CE changes over the course of training time. Most of the time, fine-tuning with LongCE loss is also a more efficient method.
@@
5 Related Work
Long-context Evaluation. Recent studies have introduced several benchmarks to evaluate the long-context performance in downstream tasks. A widely used type of benchmark is retrieval-based synthetic task, including needle-in-a-haystack (Kamradt, 2023), passkey-retrieval (Mohtashami & Jaggi, 2023) and LongEval (Li et al., 2023a). Some evaluation suites have also been gradually introduced, such as LongBench (Bai et al., 2023b), RULER (Hsieh et al., 2024), ZeroSCROLLS (Shaham et al., 2023), including document question answering, summarization, few-shot learning, code completion, and other synthetic tasks, thereby offering a more thorough evaluation of a model’s long-context abilities. To further enhance the context length of the evaluation data, InfiniteBench (Zhang et al., 2024b) has introduced evaluation data exceeding 100K tokens. In this work, we focus on understanding and addressing the limitations of perplexity at long-context training and evaluation.
Re-weighting methods in language model training. Re-weighting methods for language model training have been extensively studied, with a focus on enhancing model performance (Lin et al., 2024), improving training efficiency (Clark et al., 2022), and addressing token imbalance (Luo et al., 2023; Hu et al., 2023; Gu et al., 2020; Wang et al., 2020). Many works have also explored re-weighting through data selection techniques, addressing a wide range of challenges such as data quality (Lin et al., 2024; Li et al., 2023b), data diversity (Liu et al., 2023), and distribution matching (Li et al., 2023c; Ni et al., 2024). However, in long-context understanding, little attention was paid to the roles of individual tokens, which we find is pivotal for understanding the failure of perplexity measures. From a different perspective, a concurrent work LongRecipe (Hu et al., 2024b) also proposes reweighting tokens during training by comparing token probabilities between a base model and a trained model. Instead, we identify key tokens by computing the differences under short and long contexts, inspired by causal intervention. As a result, we only require one model during training, which saves memory footprint especially for large models.
@@ A.1 Implementation Details of LongPPL
Sliding window algorithm to improve efficiency. Since the calculation of LongPPL requires computing the LPG for each token $x_{i},i\in[n]$, it necessitates calculating the probability under short context $P_{\theta}(x_{i}|\bm{s}_{i})$ for $n$ times. Since it cannot be computed in single forward pass, this leads to massive computational overhead. To overcome this problem, we use a sliding window algorithm to improve efficiency. Specifically, we introduce a step size $d$, which is smaller than the truncation length $l$ (we set it to $d=1024$). When calculating the short-context probabilities of $x_{i}$ to $x_{i+d-1}$, we set the starting token of the context uniformly as $x_{i-l}$. Formally speaking, we have
@@
where $k\in\mathbb{N},0\leq i^{\prime}<d$. This approach allows for the calculation of the short-context probabilities of $d$ tokens in a single forward pass.
@@
Besides, in Table 1, we also implement the LongPPL with the soft influence function $I_{\rm soft}$ (Eq. (7)). In this approach, we implement an reweighting algorithm to transfer the weight between different tokenizers. Specifically, denote $\bm{w}=(w_{1},...,w_{n})$ as the LPG weight on $\bm{x}$ calculated by $P_{\theta_{0}}$. The weight of $x^{\prime}_{i}$ is defined as
@@
A.2 Implementation Details of LongPE
@@ B.1 Detailed results of LongPPL
We present the LongPPL calculated by different models in Table 4, and provide further visualization results for Mistral Large 2 and Llama-3.1-8B in Figure 8 and 9.
@@
Mixtral-8x7B-32k 2.07 2.47 1.76 3.72
FILM-7B-32k 2.51 2.95 2.04 4.54
Mistral-7B-32k 2.68 3.35 2.19 4.33
Qwen1.5-14B-128k 2.90 2.80 2.39 5.30
Qwen2-7B-128k 2.96 2.73 2.36 5.03
Phi-3-small-128k 3.01 2.87 2.46 5.48
CLEX-7B-64k 3.54 4.27 2.89 3.93
Yi-6B-200k 3.66 3.76 2.93 5.17
Yarn-7B-128k 3.67 4.82 3.09 4.20
(a) LongEval.
(b) RULER
Figure 8: LongPPL with Mistral Large 2.
(a) LongEval
(b) RULER
Figure 9: LongPPL with Llama-3.1-8B.
@@ B.2 Ablation study
LPV. In the calculation of LongPPL, we employ LPV as an assistant for our core criterion, LPG, in selecting key tokens. In Figure 10, we demonstrate the LongPPL calculated without the LPV criterion. This version of LongPPL hardly has correlation with the long-context benchmark, showing that LPV is an indispensable part for LongPPL.
Figure 10: LongPPL without LPV.
@@ B.3 Fine-grained results of LongCE
Appendix C Models

### v2

@@ 4 Experiments
In this section, we conduct real-world experiments to analyze the applicability of the proposed LongPPL and LongCE. For all the experiments, we use LongBench (Bai et al., 2023b), LongEval (Li et al., 2023a), and RULER (Hsieh et al., 2024) as the long-context benchmarks. We report the average score on LongBench, the accuracy on the subtask “lines” of LongEval, and the score on RULER. For LongBench and RULER, we restrict the prompt length to 32k tokens. For LongEval, we use 1350 lines as the prompt, which is approximately 32k tokens.
@@ 4.1 LongPPL Metric
Experimental Setup. We calculate LongPPL on the GovReport dataset (Huang et al., 2021), which consists of long sequences from government reports. We sample 50 documents with the context length up to 32k tokens. We set the hyperparameters as $\alpha=2,\beta=-2,K=4096$. We use Qwen2-72B-Instruct (Yang et al., 2024), an open-source LLM with the context length of 128k tokens, as the discriminator model $\theta_{0}$ to select the key tokens. We also consider using Llama-3.1-8B (Dubey et al., 2024) later and Mistral Large 2 (Jiang et al., 2023) in Appendix B.1.
@@
Figure 5: Correlation between the PPL-based metrics (LongPPL and PPL) on GovReport (Huang et al., 2021) and long-context benchmarks. LongPPL is calculated using Qwen2-72B-Instruct. Results of LongBench is in Figure 1(b).
(a) LongBench
(b) LongEval
(c) RULER
Figure 6: Correlation between LongPPL on GovReport and long-context benchmarks. LongPPL is calculated using Llama-3.1-8B.
LongPPL Correlates Well with Long-context Performance. In Figure 1(b) and Figure 5, we demonstrate the correlation between LongPPL and long-context benchmarks on various long-context LLMs. We observe that LongPPL exhibits a very strong negative correlation with performance on long-context tasks across different models, with pearson correlation coefficients exceeding -0.8 for all three tasks. In contrast, perplexity hardly shows a correlation with the long-context tasks. This indicates that LongPPL is sufficiently capable of measuring a model’s long-context capabilities.
Table 1: The Pearson correlation between different perplexity measures and benchmark scores, where a lower correlation is the better (since we expect a lower perplexity indicates higher benchmark scores).
@@
LongPPL-hard (default) $I$ (Equation 4) -0.96 -0.86 -0.84
LongPPL is Compatible with Small-sized Evaluator Models. To demonstrate that the effectiveness of LongPPL is not restricted by the size of the evaluator model, we additionally conduct experiments on a smaller model, Llama-3.1-8B (Dubey et al., 2024). As shown in Figure 6, the LongPPL computed using an 8B-sized model also achieves high correlation coefficients of -0.96, -0.89, and -0.90 with the three long-context benchmarks, respectively. In Appendix B.8, we have made discussion about the efficiency of LongPPL.
Hard Standard for Key Tokens is Better than Soft Re-weighting Standard. In Equation 4, we use an indicator function $I$ as the influence function. Instead, we have also tried to use the soft reweighting function $I_{\rm soft}$ used in LongCE (Equation 7) to calculate LongPPL. Its token matching strategy is detailed in Appendix A.1. In Table 1, we show that LongPPL with soft criteria has a weaker correlation with the long-context benchmarks compared to LongPPL, indicating that the soft reweighting influence function is suboptimal for LongPPL. Besides, in Appendix B.2 and B.7, we have also explored some other alternative approaches, including using the model itself as the evaluator, removing the LCL discriminative condition, and using N-grams as the key token discriminative condition. We find that all of these approaches led to worse performance.
LongPPL is not sensitive to the choice of hyperparameters of $\alpha$ and $\beta$. To investigate the impact of the two threshold hyperparameters, i.e., $\alpha$ and $\beta$ (in Equation 4), we conducted further ablation experiments. The results are presented in Table 2. Our findings reveal that when $\beta$=1, $\alpha$=1 or 2, the correlation between LongPPL and the long-context benchmarks even improves. Notably, these hyperparameters were directly reused from the motivation experiments without any further tuning. The results indicate that LongPPL’s performance is largely insensitive to the choice of hyperparameters, with the correlation coefficient remaining below -0.8 in most cases.
Table 2: The Pearson correlation between LongPPL, calculated with different hyperparameters ($\alpha$, $\beta$), and the long-context benchmarks. In most cases, the correlation coefficients remain below -0.8.
LongPPL LongBench LongEval RULER
$\alpha=2,\beta=2$ (default) -0.96 -0.86 -0.84
$\alpha=2,\beta=1$0.96 -0.92 -0.92
$\alpha=1,\beta=2$0.91 -0.73 -0.69
$\alpha=1,\beta=1$0.97 -0.88 -0.87
@@ 4.2 Fine-tune with LongCE loss
Experimental Setup. We primarily use Llama-2-7B (Touvron et al., 2023) as the base model to perform long-context finetuning. We also conduct experiments on Mistral-7B-v0.1 (Jiang et al., 2023) and Llama-2-13B. We use PG-19 (Rae et al., 2020), a book dataset sourced from a library, and Pile-arxiv (Gao et al., 2020), a dataset consisting of Arxiv papers, as the training dataset. The training sequences are organized to be the context length with 32k tokens. For the calculation of LongCE, we set $\gamma=5$ in Equation 7 and use the same sliding window approach as described in Section 4.1 to improve training efficiency. The context length of $\bm{s}_{i}$ is set to be $K=4096$. We fine-tune the base models with Entropy-aware Adjusted Base Frequency (EABF) (Zhang et al., 2024c) and Position Interpolation (PI) (Chen et al., 2023). Specifically, EABF applies a scaling mechanism to the attention and uses a higher base frequency for RoPE, while PI linearly downscales the position indices of the input tokens. These methods can significantly accelerate the convergence speed of long-context fine-tuning and have been widely adopted in many LLMs (Yang et al., 2024; Dubey et al., 2024; Chen et al., 2024a). Detailed training setups are available in Appendix A.2.
@@
LongCE Outperforms CE in Various Settings. As shown in Table 3, we present the long-context capabilities of models fine-tuned with LongCE loss and CE loss under different fine-tuning strategies and training datasets (see fine-grained results of LongBench in Appendix B.3). We also test the effectiveness of LongCE using different base models in Table 4. We find that models fine-tuned with LongCE loss consistently outperform those fine-tuned with CE loss across nearly all settings. This suggests that the LongCE loss, with its re-weighting strategy based on long-context token importance, can be applied as a plug-and-play module which can effectively improve the model’s long-context performance. To demonstrate the model’s performance when the context length is over 32K, we provide the Needle-in-a-Haystack (Kamradt, 2023) evaluation results in Appendix B.5, which leads to similar conclusions. Besides, empirical results in Appendix B.6 demonstrate that LongCE does not cause any additional loss in the model’s performance on normal-length tasks.
@@
Training Efficiency. In addition to the performance improvement brought by the LongCE loss, we also pay attention to the changes in training efficiency. In LongCE, we need an extra forward pass to calculate the probability under short context $P_{\theta}(x_{i}|\bm{s}_{i})$, which introduces additional computation costs. By using a sliding window technique (as detailed in Appendix A.1), the computational overhead of training the model with LongCE is controlled to about 80% that of training with CE loss. We visualize in Figure 7 how the long-context performance of models fine-tuned with LongCE and CE changes over the course of training time. Most of the time, fine-tuning with LongCE loss is a more efficient method. Additionally, in Appendix B.2, we find that by changing the hyperparameters of LongCE, i.e., the short context-length $K$ and the sliding window length $d$, this overhead can be further reduced to 36%, with almost no loss in model performance.
@@ 5 Conclusion
Acknowledgement
Yisen Wang was supported by National Key R&D Program of China (2022ZD0160300), National Natural Science Foundation of China (92370129, 62376010), and Beijing Nova Program (20230484344, 20240484642). Yifei Wang and Stefanie Jegelka were supported in part by the NSF AI Institute TILOS, and an Alexander von Humboldt Professorship.
@@ A.1 Implementation Details of LongPPL
Sliding window algorithm to improve efficiency. Since the calculation of LongPPL requires computing the LSD for each token $x_{i},i\in[n]$, it necessitates calculating the probability under short context $P_{\theta}(x_{i}|\bm{s}_{i})$ for $n-K$ times, where $K$ is the length of $s_{i}$. Theoretically, the computational complexity of this process is $O((n-K)K^{2})$. Since $K^{2}$ is typically larger than $n$ (e.g., when $K=4096$, $K^{2}=16$M, which is much greater than $n=32$k), this complexity far exceeds the normal $O(n^{2})$ complexity of a standard long-context forward pass. As a result, the time cost of this process is quite significant.
To make this process more efficient, we use a sliding window algorithm to improve efficiency. Specifically, we introduce a step size $d$, which is smaller than the truncation length $l$ (we set it to $d=1024$). When calculating the short-context probabilities of $x_{i}$ to $x_{i+d-1}$, we set the starting token of the context uniformly as $x_{i-l}$. Formally speaking, we have
@@
where $k\in\mathbb{N},0\leq i^{\prime}<d$. This approach allows for the calculation of the short-context probabilities of $d$ tokens in a single forward pass, resulting in a complexity of $O((N-K)K^{2}/d)$. To access a better understanding on the selection of $K$ and $d$, please refer to Appendix B.2.
@@
Besides, in Table 1, we also implement the LongPPL with the soft influence function $I_{\rm soft}$ (Eq. (7)). In this approach, we implement an reweighting algorithm to transfer the weight between different tokenizers. Specifically, denote $\bm{w}=(w_{1},...,w_{n})$ as the LSD weight on $\bm{x}$ calculated by $P_{\theta_{0}}$. The weight of $x^{\prime}_{i}$ is defined as
@@
A.2 Implementation Details of LongCE
@@ B.1 Detailed results of LongPPL
We present the LongPPL calculated by different models in Table 5, and provide further visualization results for Mistral Large 2 in Figure 8.
@@
Mixtral-8x7B-32k 2.08 2.50 1.74 3.67
FILM-7B-32k 2.49 3.17 2.03 4.47
Mistral-7B-32k 2.68 3.49 2.19 4.25
Qwen1.5-14B-128k 2.97 2.93 2.33 5.23
Qwen2-7B-128k 2.99 2.73 2.29 4.97
Phi-3-small-128k 2.98 2.86 2.41 5.42
CLEX-7B-64k 3.70 4.60 2.92 4.13
Yi-6B-200k 3.62 3.92 2.86 5.11
Yarn-7B-128k 3.67 4.88 3.10 4.17
(a) LongBench
(b) LongEval
(c) RULER
Figure 8: Correlation between LongPPL on GovReport and long-context benchmarks. LongPPL is calculated using Mistral Large 2.
@@ B.2 Ablation study
LCL. In the calculation of LongPPL, we employ LCL as an assistant for our core criterion, LSD, in selecting key tokens. In Figure 9, we demonstrate the LongPPL calculated without the LCL criterion. This version of LongPPL hardly has correlation with the long-context benchmark, showing that LCL is an indispensable part for LongPPL.
Figure 9: LongPPL without LCL.
@@
Hyperparameters of LongCE. In the computation of LongCE, several hyperparameters are utilized, including the short context window length $K$ and sliding window length $d$ used in calculating LSD. Here, we design ablation experiments to analyze the selection of these hyperparameters, as shown in Table 7. The results reveal that, on one hand, increasing $K$ or decreasing $d$ significantly improves the efficiency of LongCE (from +79% to +36%/+43%). On the other hand, under these settings, although the model’s performance on real-world tasks (LongBench) slightly decreases, it achieves substantial improvements on synthetic tasks (LongEval, RULER). This suggests that LongCE still holds potential for further efficiency enhancements.
Table 7: The performance and time cost of LongCE on long-context benchmarks under different hyperparameter settings of $K$ and $d$. For the time cost, we report the wall-clock time for training 200 steps.
Total training time / h LongBench LongEval RULER
Training steps 200 50 100 200 50 100 200 50 100 200
Setting A (PG-19 dataset with EABF)
CE 7.0 24.5 26.6 26.9 16.0 24.0 24.0 34.5 38.6 42.7
LongCE ($K=4$k, $d=1$k, default) 12.5 (+79%) 26.0 27.2 28.2 24.0 46.0 46.0 43.1 48.3 49.7
LongCE ($K=1$k, $d=1$k) 10.0 (+43%) 25.3 25.8 26.9 20.0 48.0 48.0 45.6 51.1 55.9
LongCE ($K=4$k, $d=4$k) 9.5 (+36%) 25.4 25.8 25.8 28.0 56.0 56.0 42.5 48.0 51.2
LongCE ($K=4$k, $d=512$) 17.5 (+150%) 25.4 25.8 27.3 26.0 48.0 60.0 42.4 50.1 54.4
@@ B.3 Fine-grained results of LongCE
B.4 Detailed results of the experiments in section 2.1
In Table 9, we present the detailed results from the experiments in Figure 2(b) and 2(c).
Table 9: Detailed results of experiments in Figure 2, including the accuracy on LongEval, and perplexity tested on answer and non-answers tokens, respectively.
Prompt Length 2k 3k 4k 5k 7k 9k 11k 13k 15k 17k 19k 21k 23k 25k 28k
Yi-6B-200K
LongEval accuracy / % 100.0 94.0 84.0 76.0 76.0 64.0 68.0 54.0 60.0 58.0 46.0 44.0 50.0 52.0 48.0
PPL (answer tokens) 1.49 1.47 1.59 1.64 1.91 2.00 1.98 2.29 2.28 2.15 2.39 2.11 2.23 2.32 2.08
PPL (non-answer tokens) 2.15 2.17 2.12 2.18 2.18 2.20 2.27 2.25 2.25 2.23 2.23 2.21 2.22 2.25 2.24
CLEX-7B-64K
LongEval accuracy / % 82.0 34.0 84.0 82.0 58.0 62.0 58.0 56.0 50.0 44.0 46.0 24.0 22.0 28.0 24.0
PPL (answer tokens) 1.31 2.33 1.23 1.33 1.47 1.43 1.51 1.54 1.63 1.78 1.89 2.23 2.50 2.61 2.59
PPL (non-answer tokens) 2.22 2.31 2.17 2.18 2.10 2.16 2.17 2.14 2.14 2.15 2.15 2.18 2.20 2.24 2.24
B.5 Needle-in-a-haystack results
In this section, we conduct the standard Needle-in-a-Haystack (NIAH) evaluation to evaluate models’ long-context capability when context lengths is greater than 32K.
We first test the models obtained in the main text, which are fine-tuned on 32K-length texts. As shown in figure 10, LongCE achieves a score of 10 on 5 out of 6 questions at the 40K length and 2 out of 6 questions at the 48K length, outperforming CE, which achieves a score of 10 on 2 out of 6 and 0 out of 6 questions, respectively. Therefore, LongCE demonstrates a longer effective context length.
Additionally, to demonstrate the generalization ability of LongCE on longer context lengths, we extend the context window of both models by increasing their RoPE base from 500K to 2M. The corresponding NIAH results are shown in Figure 11. The results show that model finetuned with LongCE answers all questions correctly at the 64K length and achieves a score of 10 on 32 sequences with lengths of $\geq$32K, while CE only achieves this on 26 sequences. This indicates that LongCE can generalize well at longer lengths.
(a) Model finetuned with CE.
(b) Model finetuned with LongCE.
Figure 10: Needle-in-a-haystack results of models trained with PG-19 datasets & EABF for 200steps.
(a) Model finetuned with CE.
(b) Model finetuned with LongCE.
Figure 11: Needle-in-a-haystack results of models trained with PG-19 datasets & EABF for 200steps. We increase the RoPE base from 500k to 2M after finetuning.
B.6 LongCE’s performance on non-long-context language tasks
In this section, we experimentally investigate whether LongCE will adversely impact non-long-context capabilities. In Table 10, we present the model performance on 6 common language tasks, i.e., MMLU (Hendrycks et al., 2021), ARC-Challenge (Clark et al., 2018), RACE (Lai et al., 2017), BigBench Hard (Suzgun et al., 2023), TruthfulQA (Lin et al., 2022), and CommonsenseQA (Talmor et al., 2019). The results show that for non-long-context tasks, the performance of the model trained with LongCE is nearly identical to that of the model trained with CE, indicating that the long-context-specific characteristics of LongCE do not negatively affect the model’s performance on tasks involving normal-length context compared to the baseline.
Table 10: The performance of models fine-tuned with CE and LongCE on non-long-context tasks. The models are finetuned with 200 steps under the setting A in Table 3.
Models MMLU ARC-C RACE BBH TruthfulQA CommonsenseQA Avg.
Llama-2-7B 41.8 43.3 39.5 39.4 34.5 32.9 38.6
   +CE (baseline) 40.8 42.8 40.3 36.4 29.3 31.5 36.9
   +LongCE (ours) 39.9 43.9 39.3 37.5 30.0 30.8 36.9
B.7 Substituting key tokens with re-occurred N-gram
In this section, we examine whether LongPPL works by retrieving the frequent N-gram in the context, as concerned in recent works (Sun et al., 2021; Arora et al., 2024). We calculate perplexity solely on the re-occurred N-gram (word-level, $N>2$) in the inputs, and present the correlation coefficients with the benchmarks in Table 11.
Table 11: The correlation coefficients between PPL calculated on re-occurred N-gram, and the benchmarks.
LongBench LongEval RULER
PPL -0.11 0.24 0.27
PPL (N-gram) -0.41 -0.10 -0.05
LongPPL -0.96 -0.86 -0.84
The results show that PPL on re-occurred N-grams has much weaker correlation with model’s long-context capabilities. This indicates that LongPPL’s powerful ability to capture long-context-related information cannot be simply explained by N-grams.
B.8 Time consumption of LongPPL
In Table 12, we test the time cost of LongPPL. It can be observed that the time cost of calculating LongPPL using the 8B model as the evaluator is approximately 3$\sim$4 times that of calculating PPL, while the overhead for using the 72B model is much higher.
Although the computational overhead of LongPPL is non-negligible, we believe that such a computational cost will not have a substantial impact on the practicality of LongPPL. On the one hand, if users employ LongPPL as a benchmark, key tokens can be calculated offline, resulting in no online computation overhead. On the other hand, if LongPPL is used as an evaluation metric during training, its computational overhead is negligible compared to the overall training cost (as evaluation steps are typically sparse during training).
Table 12: The time consumption of LongPPL. The values in the table represent the average seconds required per sequence.
PPL LongPPL (Llama-3.1-8B) LongPPL (Qwen2-72B-Instruct)
Mistral-7B 2.8 11.3 (+8.5, +304%) 56.4 (+53.6, +2014%)
Mixtral-8x7B (47B) 4.2 13.5 (+9.3, +221%) 58.4 (+54.2, +1390%)
Appendix C Related Work
Long-context Evaluation. Recent studies have introduced several benchmarks to evaluate the long-context performance in downstream tasks. A widely used type of benchmark is retrieval-based synthetic task, including needle-in-a-haystack (Kamradt, 2023), passkey-retrieval (Mohtashami & Jaggi, 2023) and LongEval (Li et al., 2023a). Some evaluation suites have also been gradually introduced, such as LongBench (Bai et al., 2023b), RULER (Hsieh et al., 2024), ZeroSCROLLS (Shaham et al., 2023), including document question answering, summarization, few-shot learning, code completion, and other synthetic tasks, thereby offering a more thorough evaluation of a model’s long-context abilities. To further enhance the context length of the evaluation data, InfiniteBench (Zhang et al., 2024b) has introduced evaluation data exceeding 100K tokens. In this paper, we analyze the correlation between the Perplexity metric and specific evaluation tasks and propose an alternative LongPPL metric, which can better align the model’s long-context performance on downstream tasks.
Re-weighting methods in language model training. Re-weighting methods for language model training have been extensively studied, with a focus on enhancing model performance (Lin et al., 2024), improving training efficiency (Clark et al., 2022), and addressing token imbalance (Luo et al., 2023; Hu et al., 2023; Gu et al., 2020; Wang et al., 2020). Many works have also explored re-weighting through data selection techniques, addressing a wide range of challenges such as data quality (Li et al., 2023b), data diversity (Liu et al., 2023), and distribution matching (Li et al., 2023c; Ni et al., 2024). However, few of these works focus on re-weighting tokens to enhance a model’s long-context performance. The most recent and closely related work to ours is LongRecipe (Hu et al., 2024b), which re-weights tokens based on distribution shifts in model predictions during training. This approach does not capture the essential characteristics of key tokens. In contrast, our method directly re-weights tokens according to their dependence on long-context information, providing a more fundamental and targeted solution.
Appendix D Models